OpenStudy (anonymous):

3squareroot2-2squareroot3 over 3squareroot2+2squareroot3 Rationalize the denominator and simplify.

OpenStudy (anonymous):

\[3\sqrt{2}-2\sqrt{3}\over3\sqrt{2}+2\sqrt{3}\]

OpenStudy (anonymous):

\[\frac{ a-b }{ a+b }*\frac{ a-b }{ a-b }\] change the sign of denominator and multiply both in numerator and denominator.

OpenStudy (anonymous):

have you followed..? show me your work.

OpenStudy (anonymous):

Sorry, just got back on.. And im lost bc im not timesing it

OpenStudy (anonymous):

\[\frac{ 3\sqrt{2}-2\sqrt{3} }{3\sqrt{2}+2\sqrt{3} }*\frac{ 3\sqrt{2}-2\sqrt{3} }{ 3\sqrt{2}-2\sqrt{3} }\] \[for denominator \left( a+b \right)\left( a-b \right)=a ^{2}-b ^{2}\] \[for numerator \left( a-b \right)*\left( a-b \right)=a ^{2}-2*a*b+b ^{2}\]

OpenStudy (anonymous):

But what do you do with the square roots??

OpenStudy (anonymous):

\[example \left( 5\sqrt{6} \right)^{2}=5^{2}*6=25*6=150\]

OpenStudy (anonymous):

\[5\sqrt{6}*3\sqrt{7}=5*3 \sqrt{6*7}=15\sqrt{42}\]

OpenStudy (anonymous):

So I would take \[3\sqrt{2} (5\sqrt{6})?\]

OpenStudy (anonymous):

these are examples only. you stick to your question.

OpenStudy (anonymous):

Yeah now I'm all confused.

OpenStudy (anonymous):

1 have told how to square a term with square root In the second example i have told how to multiply two terms with squareroots.

OpenStudy (anonymous):

Well I'm sorry, I'm a dumbass and can't understand this worth pellet.

OpenStudy (anonymous):

\[you have tofind thesquare of 3\sqrt{2} and 2\sqrt{3}\]

OpenStudy (anonymous):

Okay, now how do I do that?

OpenStudy (anonymous):

\[\left( 3\sqrt{2} \right)^{2}=3^{2}*2=9*2=18\] \[similarly \left( 2\sqrt{3} \right)^{2}=2^{2}*3=4*3=12\] \[3 \sqrt{2}*2\sqrt{3}=3*2\sqrt{2*3}=6\sqrt{6}\] i think now you can complete.

OpenStudy (anonymous):

Wouldn't that be \[3\sqrt{2}?\]

OpenStudy (anonymous):

no, terms under square root are multiplied. and terms out square root are multiplied separately.